Martina leaves home at 9 A.M., bicycling at a rate of 24mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by checkley71(8403)

24*T=48(T-2) OR 24T=48T-96 OR 96=24T OR T=96/24 OR T=4 OR JOHN CATCHES MARTINA @ 1:00PM.